OpenStudy (elleblythe):

How to find the exact value of csc(5pi/12)

OpenStudy (gorv):

5pi/12 can u break it into 2 part

OpenStudy (gorv):

if we take it in degree

OpenStudy (gorv):

\[5*\frac{ 180 }{ 12 }=5*15\]

OpenStudy (gorv):

75 degree

OpenStudy (gorv):

we can write it as 45 +30 degree

OpenStudy (gorv):

@elleblythe

OpenStudy (gorv):

\[\csc(\frac{ \pi }{ 4 }+\frac{ \pi }{ 6 })\]

OpenStudy (anonymous):

csc(2pi/12+3pi/12)=csc(pi/6+pi/4) =1/(sin(pi/6+pi/4)) =1/(sin(pi/6).cos(pi/4)+cos(pi/6).sin(pi/4))